# Multi-Agent Obstacle Avoidance using Velocity Obstacles and Control Barrier Functions

### v2

@@ V Experiments
Figure 3: Visualization of the resulting trajectories in the baseline comparison study. The arrows in each corresponding color describes the direction of motion of each agent. Our method (left-most column) remains stable and computes smooth trajectories for all evaluated scenarios. The hVO method encountered infeasibility (no available safe velocities) in the scenarios with $8$ and $12$ agents. The aspect ratio is equal for both dimensions and the scale is given by the bottom left reference.
@@ V-A Setup
Parameter 2nd Integrator Car Dynamics
@@ V-B1 Setup
As baselines, we used Velocity Obstacles (VO), Reciprocal Velocity Obstacles (RVO), Collision Cone CBF (hVO), and Optimal Velocity selection using Velocity Obstacle (OVVO). For implementation, we followed the methodology outlined in [15, 2], while focusing on the 2nd-order integrator dynamics. To handle the fact that the other approaches had no explicit motion model, with control $\mathbf{u}$, we did as follows. At each timestep, we define and sample a set of admissible velocities based on the maximum allowable acceleration and timestep. For VO and RVO, we select the one that minimizes (analogous to (16)) the penalty
$J_{\mathrm{VO},\mathrm{RVO}}=\frac{1}{TTC}+||\mathbf{v}-\mathbf{v}_{\mathrm{des}}||,$ (32)
where TTC is the minimum TTC among obstacles. For OVVO, the penalty is expressed as
@@
where $\frac{k_{tp}}{k_{vd}}$ balances deviating from the VO cone and following the desired control input, and $c_{1}$, $c_{2}$ are constants. $d_{v}$ and $t_{p}$ represent the clearance and pass time, closely related to a separation between obstacles and time-to-collision respectively (we refer the reader to [2] for further details). We recover the hVO formulation by setting $k_{cone}$ and the slack variables ($\lambda_{i}$) to 0 while removing the safety-critical constraints.
The time-to-collision (TTC) is computed assuming constant velocity, solving for the time $t$ such that
@@
where $\mathbf{v}_{A}^{\prime}$ is the desired velocity for agent $A$. The selected velocity is divided by the timestep to obtain the control input. All models’ parameters are manually tuned with minimal effort to evaluate qualitative performance. No optimization is applied. The control scheme is PD-based, using position error as the input to guide the agents toward their goals.
We evaluate our method in a circular formation scenario to compare its performance against the baselines. Agents are equidistantly positioned on a circle with zero initial velocity, and their goal is to reach the diametrically opposite point. This naturally creates conflicting trajectories, leading to potential collisions. As the system is deterministic, we add noise to the initial position to compute statistics. We examine cases with 2, 4, 8, and 12 agents, all governed by 2nd-order integrator dynamics. Each scenario is run 10 times and statistics are reported.
@@ V-B2 Results
Results are summarized in Table II and visualized in Figure 3. Qualitatively, our method produces smoother trajectories compared to the baselines. While all methods perform well with few agents, increasing the number of agents resulted in greater challenges, particularly for VO, hVO, and OVVO, leading to numerous collisions or low success rates. RVO performs better than the others, but still could not completely avoid collisions, as seen in the 12-agent case. In contrast, our approach consistently yielded zero collisions, as predicted by the safety guarantees.
In terms of success rate, hVO fails at 8 and 12-agent scenarios. At the start of the simulations, all agents are in a safe state with 0 velocity. This leads them to choose a velocity towards their goals. Thus, at the next timestep all of them are facing each other with colliding velocities, and since the scenario is crowded and acceleration is constrained, no feasible solution is found to the problem without relaxing the constraints. In terms of completion time, RVO performed best, with our approach second, but never needing more that 10% extra time to finish.
@@
Thus, while our method is similar to RVO in overall performance, it provides safety guarantees. Moreover, the smoother trajectories might provide less wear and tear in hardware over time.
RVO, while faster, assumes a degree of cooperation among agents, whereas our method is more general and can be employed in decentralized systems. Since the constraints are linear, we get a Quadratic Programming problem (QP), keeping computational complexity unchanged—an important advantage for scaling to larger problems.
@@ V-C Car Dynamics Example
We extend the experiments to car-like dynamics, see Equation (15), maintaining the same methodology. Two scenarios are evaluated: one with 4 agents and another with 8 agents, where the agents are tasked to reach the diametrically-opposite location. The results are consistent with the 2nd-order integrator dynamics. As can be seen in Figure 4, the resulting trajectories are smooth, and the non-holonomic agents successfully adapted to each other.
Videos of the presented experiments can be found at https://youtu.be/Ox8v2s17gLw.
@@ VI Conclusion
In this paper we have shown how to combine CBFs with VO in a way that guarantees safety, makes use of the potential efficiency provided by the VO by including it in the objective of the optimization, overcoming the downside of being overly conservative that results from considering VOs as hard constraints. The simulation comparisons show that the proposed approach goes beyond the existing ones in terms of path smoothness, success rates and collision rates.

### v3

@@ V Experiments
Code [^1] https://github.com/KTH-RPL/MA-CBF-VO and videos [^2] https://youtu.be/Ox8v2s17gLw of the experiments are available online.
Figure 3: Visualization of the resulting trajectories in the baseline comparison study. The arrows in each corresponding color describe the direction of motion of each agent. Our method (left-most column) remains stable and computes smooth trajectories for all evaluated scenarios. The hVO method encountered infeasibility (no available safe velocities) in the scenarios with $8$ and $12$ agents. The aspect ratio is equal for both dimensions and the scale is given by the bottom left reference.
@@ V-A Setup
Parameter Second-order Int. Car Dynamics
@@ V-B1 Setup
As baselines, we used Velocity Obstacles (VO), Reciprocal Velocity Obstacles (RVO), Collision Cone CBF (hVO), and Optimal Velocity selection using Velocity Obstacle (OVVO). For implementation, we followed the methodology outlined in [15, 2], while focusing on the second-order integrator dynamics. To handle the fact that the other approaches had no explicit motion model, with control $\mathbf{u}$, we did as follows. At each timestep, we define and sample a set of admissible velocities based on the maximum allowable acceleration and timestep. For VO and RVO, we select the one that minimizes (analogous to (16)) the penalty
$J_{\mathrm{VO},\mathrm{RVO}}=\frac{1}{T_{\mathrm{col}}}+||\mathbf{v}-\mathbf{v}_{\mathrm{des}}||,$ (32)
where $T_{\mathrm{col}}$ is the minimum time-to-collision among obstacles. For OVVO, the penalty is expressed as
@@
where $\nicefrac{{k_{tp}}}{{k_{vd}}}$ balances deviating from the VO cone and following the desired control input, and $c_{1}$, $c_{2}$ are constants. $d_{v}$ and $t_{p}$ represent the clearance and pass time, closely related to a separation between obstacles and time-to-collision respectively (we refer the reader to [2] for further details). We recover the hVO formulation by setting $k_{cone}$ and the slack variables ($\lambda_{i}$) to 0 while removing the safety-critical constraints.
The time-to-collision ($T_{\mathrm{col}}$) is computed assuming constant velocity, solving for the time $t$ such that
@@
where $\mathbf{v}_{A}^{\prime}$ is the desired velocity for agent $A$. The selected velocity is divided by the timestep to obtain the control input. All parameters of the models are manually tuned with minimal effort to evaluate qualitative performance. No optimization is applied. The control scheme is PD-based, using position error as the input to guide the agents toward their goals.
We evaluate our method in a circular formation scenario to compare its performance against the baselines. Agents are equidistantly positioned on a circle with zero initial velocity, and their goal is to reach the diametrically opposite point. This naturally creates conflicting trajectories, leading to potential collisions. As the system is deterministic, we add noise to the initial position to compute statistics. We examine cases with 2, 4, 8, and 12 agents, all governed by second-order integrator dynamics. Each scenario is run 10 times and statistics are reported.
@@ V-B2 Results
Results are summarized in Table II and visualized in Figure 3. Qualitatively, our method produces smoother trajectories compared to the baselines. While all methods perform well with few agents, increasing the number of agents resulted in greater challenges, particularly for VO, hVO, and OVVO, leading to numerous collisions or low success rates. RVO performs relatively well, but still could not completely avoid collisions, as seen in the 12-agent case. In contrast, our approach consistently yielded zero collisions, as predicted by the safety guarantees.
In terms of success rate, hVO fails at 8 and 12-agent scenarios. At the start of the simulations, all agents are in a safe state with zero velocity. This leads them to choose a velocity towards their goals. Thus, at the next timestep, all of them are facing each other with colliding velocities, and since the scenario is crowded and acceleration is constrained, no feasible solution is found to the problem without relaxing the constraints. In terms of completion time, RVO performed best, with our approach second, but never needing more than 10% extra time to finish.
@@
Thus, while our method is similar to RVO in overall performance, it provides safety guarantees. Moreover, the smoother trajectories might provide less wear and tear in hardware over time. RVO assumes a degree of cooperation among agents, whereas our method is more general and can be employed in decentralized systems. Since the constraints are linear, we get a Quadratic Programming problem (QP), keeping computational complexity unchanged—an important advantage for scaling to larger problems.
@@ V-C Car Dynamics Example
We extend the experiments to car-like dynamics, see Equation (15), maintaining the same methodology. Two scenarios are evaluated: one with 4 agents and another with 8 agents, where the agents are tasked to reach the diametrically opposite location. The results are consistent with the second-order integrator dynamics. As can be seen in Figure 4, the resulting trajectories are smooth, and the non-holonomic agents successfully adapted to each other.
@@ VI Conclusion
In this paper we have shown how to combine CBFs with VO in a way that guarantees safety, and leverages the efficiency provided by the VO by including it in the objective of the optimization, overcoming the downside of being overly conservative that results from considering VOs as hard constraints. The simulation comparisons show that the proposed approach outperforms the alternatives in terms of path smoothness, success rates, and collision rates.